OpenStudy (anonymous):

Find the slope of the graph of the function at (-1, 3): h(x)=sqrt(x+10)

OpenStudy (amistre64):

find the derivative

OpenStudy (anonymous):

I found it to be \[(\sqrt{9-h}-3)/h\] When I plug in the value of h, I still get 0/0

OpenStudy (amistre64):

... you still using first principles? i take it you havent learned the derivative rules yet then

OpenStudy (anonymous):

Yeah, we are just starting into this stuff

OpenStudy (amistre64):

well, i spose i can run the a first principles; you most likely missed a conjugate

OpenStudy (amistre64):

\[\frac{\sqrt{x+h}-\sqrt{x}}{h}\] \[\frac{\sqrt{x+h}-\sqrt{x}}{h}*\frac{\sqrt{x+h}+\sqrt{x}}{\sqrt{x+h}+\sqrt{x}}=\frac{x+h-x}{h(\sqrt{x+h}+\sqrt{x})}\] \[\frac{h}{h(\sqrt{x+h}+\sqrt{x})}\] \[\frac{1}{\sqrt{x+h}+\sqrt{x}};when\ h=0\] \[\frac{1}{\sqrt{x}+\sqrt{x}}=\frac{1}{2\sqrt{x}}\]

OpenStudy (amistre64):

just replace x with x+10 and its good

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

yep